OpenStudy (anonymous):

The area of a circle is 45.6 in². The length of the radius is _____in. 3.4 in 3.8 in 4.2 in 4.6 in

OpenStudy (anonymous):

A = pi*r^2

OpenStudy (anonymous):

\[area= \pi r^{2}\] \[\sqrt{45.6/\pi} = r\]

OpenStudy (anonymous):

3.8?

OpenStudy (lalaly):

yup:D

OpenStudy (anonymous):

yaaaaaaaaaaaaaay